### v1

# Transforming LLMs into Cross-modal and Cross-lingual Retrieval Systems

Abstract
Large language models (LLMs) are trained on text-only data that go far beyond the languages with paired speech and text data. At the same time, Dual Encoder (DE) based retrieval systems project queries and documents into the same embedding space and have demonstrated their success in retrieval and bi-text mining. To match speech and text in many languages, we propose using LLMs to initialize multi-modal DE retrieval systems. Unlike traditional methods, our system doesn’t require speech data during LLM pre-training and can exploit LLM’s multilingual text understanding capabilities to match speech and text in languages unseen during retrieval training. Our multi-modal LLM-based retrieval system is capable of matching speech and text in 102 languages despite only training on 21 languages. Our system outperforms previous systems trained explicitly on all 102 languages. We achieve a 10% absolute improvement in Recall@1 averaged across these languages. Additionally, our model demonstrates cross-lingual speech and text matching, which is further enhanced by readily available machine translation data.
Transforming LLMs into Cross-modal and Cross-lingual Retrieval Systems
Frank Palma Gomez1 Ramon Sanabria2 Yun-hsuan Sung4 Daniel Cer4 Siddharth Dalmia3‡ Gustavo Hernandez Abrego4‡ 1Boston University 2The University of Edinburgh 3Google DeepMind 4Google Research fpg@bu.comthanks: Work done by Frank and Ramon during their internship in Google Research and Google DeepMind respectively.thanks: Equal Advising Contributions.
1 Introduction
LLMs have demonstrated their effectiveness in modelling textual sequences to tackle various downstream tasks Brown et al. (2020); Hoffmann et al. (2022); Chowdhery et al. (2023). This effectiveness has led to the development of powerful LLMs capable of modelling text in a wide range of languages. The abundance of textual data in different languages across the internet has fueled the progress of multi-lingual models Johnson et al. (2017); Xue et al. (2020); Siddhant et al. (2022). On the other hand, speech technologies are prevalent in smartphones and personal assistants, but their language availability is relatively limited compared to the languages that LLMs support Baevski et al. (2020); Radford et al. (2023).
Figure 1: Our dual encoder architecture and training pipeline. We expand the embedding layer of our backbone LLM to support the additional discretized speech tokens, that are extracted from a pre-trained speech encoder. At the same time, we tokenize the corresponding transcripts with the LLM tokenizer. We encode the speech tokens and transcripts separately and train the model with a contrastive loss over the dot product between speech and transcript embeddings.
Various efforts have explored solutions to the speech-text data scarcity problem Duquenne et al. (2021); Ardila et al. (2019); Wang et al. (2020). Works such as SpeechMatrix Duquenne et al. (2022) use separate speech and text encoders to mine semantically similar utterances that are neighbors in an embedding space. However, these approaches are limiting because they require speech and text encoders that have aligned representation spaces.
We posit that we can retrieve speech and text utterances by aligning both modalities within the embedding space built from a single pre-trained LLM. We take inspiration from previous works that use pre-trained LLMs to perform automatic speech recognition (ASR) and automatic speech translation (AST) Rubenstein et al. (2023); Wang et al. (2023); Hassid et al. (2023). Our intuition is that we can perform the speech and text alignment leveraging the capabilities of text-only LLMs without requiring two separate models.
In this paper, we propose converting LLMs into speech and text DE retrieval systems without requiring speech pre-training and outperform previous methods with significantly less data. By discretizing speech into acoustic units Hsu et al. (2021), we extend our LLMs embedding layer and treat the acoustic units as ordinary text tokens. Consequently, we transform our LLM into a retrieval system via a contrastive loss allowing us to match speech and text utterances in various languages. Our contributions are the following:
  1. 1.
    We build a speech-to-text symmetric DE from a pre-trained LLM. We show that our retrieval system is effective matching speech and text in 102 languages of FLEURS Conneau et al. (2023) despite only training on 21 languages.
  2. 2.
    We show that our model exhibits cross-lingual speech and text matching without training on this type of data. At the same time, we find that cross-lingual speech and text matching is furthered improved by training on readily available machine translation data.
2 Method
We train a transformer-based DE model that encodes speech and text given a dataset, $\emph{D}=\{(x_{i},y_{i})\}$, where $x_{i}$ is a speech utterance and $y_{i}$ is its transcription. We denote the speech and text embeddings as $\bm{x_{i}}=E(x_{i})$ and $\bm{y_{i}}=E(y_{i})$, respectively, where $E$ is a transformer-based DE that encodes speech and text.
2.1 Generating Audio Tokens
We convert raw speech into discrete tokens using the process in Lakhotia et al. (2021); Borsos et al. (2023). The process converts a speech query $x_{i}$ into an embedding using a pre-trained speech encoder. The output embedding is then discretized into a set of tokens using k-means clustering. We refer to the resulting tokens as audio tokens. We use the 2B variant of the Universal Speech Model (USM) encoder Zhang et al. (2023) as the speech encoder and take the middle layer as the embedding for $x_{i}$. Additionally, we generate audio tokens at 25Hz using k-means clustering, resulting in a set of 1024 possible audio tokens. We will refer to this as our audio token vocabulary.
2.2 Supporting Text and Audio Tokens
To support text and audio tokens in our LLM, we follow the formulation of Rubenstein et al. (2023). We extend the embedding layer of a transformer decoder by $a$ tokens, where $a$ represents the size of our audio token vocabulary. This modification leads to an embedding layer with size $(t+a)\times m$, where $t$ is the number of tokens in the text vocabulary and $m$ is the dimensions of the embedding vectors. In our implementation, the first $t$ tokens represent text and the remaining $a$ tokens are reserved for audio. We initialize the embeddings layer from scratch when training our model.
3 Data and Tasks
Appendix 5 details our training and evaluation datasets along with the number of languages in each dataset, the split we used, and the size of each dataset. We focus on the following retrieval tasks:
Speech-to-Text Retrieval (S2T)
involves retrieving the corresponding transcription from a database given a speech sample. In S2T, we train on CoVoST-2 Wang et al. (2021) speech utterances and their transcriptions. CoVoST-2 is a large multilingual speech corpus derived from Wikipedia expanding over 21 languages and provides translation to and from English. We use FLEURS Conneau et al. (2023) to evaluate S2T performance on 102 languages. FLEURS is an $n$-way parallel dataset containing speech utterances from FLoRES-101 Goyal et al. (2021) human translations. To evaluate S2T, we report recall at 1 ($R@1$) rates for retrieving the correct transcription for every speech sample and word error rate (WER).
Speech-to-Text Translation Retrieval (S2TT)
attempts to retrieve the corresponding text translation of a speech sample. We use S2TT to measure the cross-lingual capabilities of our multi-modal DE retrieval system. We evaluate this capability zero-shot on X $\to$ En S2TT data of FLUERS and explore if we can further improve this capability by training on readily-available machine translation data from WikiMatrix Schwenk et al. (2019). We pick French, German, Dutch, and Polish to English that are common across WikiMatrix and FLEURS and further discuss the amount of machine translation data used in Appendix 5. For S2TT, we report 4-gram corpusBLEU Post (2018).
4 Model
Figure 1 shows an illustration of our model. We initialize our dual encoder from PaLM 2 XXS Google et al. (2023) and append a linear projection layer after pooling the outputs along the sequence length dimension. The embedding and linear projection layers are initialized randomly. After initializing our model from PaLM 2, we use a contrastive loss Hadsell et al. (2006). Appendix A.1 includes more details on our training setup.
5 Experiments
R@1 $\uparrow$ WER $\downarrow$
PaLM 2 XXS 86.15 13.85
mSLAM 76.9 14.6
Table 1: PaLM 2 results for $R@1$ and WER compared against the mSLAM on 102 languages from FLEURS for speech-to-text retrieval (S2T).
We train our DE model to perform S2T, where the task is to retrieve the corresponding transcription given a speech sample. We train on the 21 languages from CoVoST-2 and evaluate our model using the S2T portion of FLEURS in 102 languages.
5.1 Speech-to-Text Retrieval
Table 1 shows the average $R@1$ and WER for S2T for 102 languages from FLEURS. We compare against the mSLAM model from Conneau et al. (2023), a model trained on 426k hours of S2T data in 51 languages and fine-tuned on FLEURS training data. Our model significantly outperforms the mSLAM baseline in $R@1$ and WER metrics despite being trained with only 1/10 of the data and having been initialized from a text-only LLM. More importantly, our model was only trained on the 21 languages in CoVoST-2 and never fine-tuned on the FLEURS training data.
5.1.1 Seen-Unseen Breakdown
Figure 2: $R@1$ transcription retrieval for seen and unseen languages in the training set.
In Figure 2 we break down the $R@1$ scores based on seen and unseen languages during training. We find that our model performs best on the 20 languages that are within the training and evaluation data, but still perform significantly well on the remaining 82 unseen languages. We hypothesize this is due to the vast textual multilingual data our backbone LLM has seen during pre-training.
5.1.2 Language Group Breakdown
R@1 $\uparrow$
Language Group (#) PaLM 2 (%) mSLAM (%)
Afro-Asiatic (7) 90.82 73.67
Atlantic-Congo (15) 79.47 86.77
Austro-Asiatic (3) 41.71 47.90
Austronesian (6) 85.74 75.50
Dravidian (4) 90.46 65.70
Indo-European (51) 92.38 84.62
Japonic (1) 63.23 5.80
Kartvelian (1) 74.57 70.50
Koreanic (1) 45.81 5.20
Kra-Dai (1) 37.12 3.20
Mongolic (1) 98.10 70.70
Nilo-Saharan (1) 94.53 91.00
Sino-Tibetan (3) 61.91 3.40
Turkic (4) 91.89 81.28
Uralic (3) 93.38 91.40
Table 2: FLEURS S2T (R@1) performance broken down by language groups. Bold represents better performance. Numbers in parenthesis represent the number of languages within the language group.
Table 2 shows the $R@1$ language group breakdown for S2T on FLEURS. We find that although we only trained on 21 languages, our model significantly outperforms mSLAM in 13 of the 15 language groups. These results are consistent with the experiments in Hassid et al. (2023) which explore the effect of initializing speech language models from pre-trained LLMs.
5.2 Evaluating on Cross-Modal and Cross-Lingual Tasks
Figure 3: BLEU scores for FLEURS zero-shot S2TT when training on Transcripts or Transcripts + Translations. Combining transcripts and translation data improves zero-shot S2TT retrieval.
We evaluate on S2TT to gauge the cross-modal and cross-lingual capabilities of our model. We show that we can further improve S2TT by simply training on a mixture of S2T and translation data without using any S2TT training data.
5.2.1 Zero-Shot S2TT
Given the multi-lingual capabilities of our backbone language model, we explore if these capabilities are transferred after training our model contrastively on the S2T task. We hypothesize that our model should showcase cross-lingual and cross-modal capabilities due to the cross-modal training task and the cross-lingual capabilities of the backbone LLM. We evaluate S2TT in a zero-shot setting to assess our model’s performance retrieving English translations given a speech sample in another language. Using the FLEURS S2TT portion, we evaluate S2TT X $\to$ En in 4 languages: German, Polish, French, and Dutch.
Figure 3 shows BLEU S2TT performance using S2T CoVoST-2 in 21 languages. We call this setup Transcripts in Figure 3. Our results demonstrate that even when only training our model on speech and transcriptions, we can achieve some zero-shot S2TT performance and We find that S2TT BLEU scores are considerably higher for languages present S2T training data. For example, Polish was not in the S2T training therefore its BLEU scores are the lowest.
5.2.2 Improving S2TT with MT Data
To further improve our model’s cross-lingual performance, we add readily available translation data from Schwenk et al. (2019) to improve S2TT. For each batch, we combine 25% translation and 75% S2T data. Figure 3 shows comparison of only training on S2T (Transcripts) and combining S2T and translation data ( Transcriptions + Translations). We find that combining S2T and translation data significantly improves the S2TT BLEU scores in all 4 languages without training on S2TT data. This finding demonstrates that we can improve our models cross-lingual performance with highly accessible translation data without needing scarce and often expensive speech-to-text translation training data.
6 Related Work
The success of pre-trained LLMs have motivated the application of these models in different modalities. Lakhotia et al. (2021) transformed speech into pseudo-text units to introduce the task of generative spoken language modeling. Borsos et al. (2023) introduced a framework to generate audio with long-term consistency. Consequently, Hassid et al. (2023) showed that SpeechLMs benefit from being initialized from pre-train LLMs while Rubenstein et al. (2023) demonstrated that pre-trained LLMs can be adapted to various tasks that required text and speech understanding.
On the other hand, several works aim to build joint speech and text representations. Chung et al. (2021) introduced w2v-bert which combines masked language modeling and contrastive learning to create speech representations. Bapna et al. (2022) jointly pre-trains on speech and text from unsupervised speech and text data. Recently, Duquenne et al. (2023) employed separate speech and text encoders to generate embeddings in over 200 languages. Nevertheless, there is still a lack of understanding of whether joint speech and text representations can be built from a single encoder. We fill this gap by using pre-trained LLMs to jointly train on speech samples and their transcriptions to show that our approach is capable of speech-text matching in 102 languages.
7 Conclusion
We present an effective approach to developing a speech-to-text DE from a text-only LLM. Our findings suggest that by using a text-only LLM as a backbone model, we can drastically outperform previous approaches using considerably less speech-to-text training data. Additionally, we find that we can improve zero-shot speech translation by simply combining readily available translation and S2T data. We showcase our findings in 102 languages for S2T and 4 languages in S2TT; opening up the possibility of using speech-to-text DE’s in different cross-model and cross-lingual settings.
8 Acknowledgements
We would like to thank Shankar Kumar and Ankur Bapna for the valuable feedback on the draft of the paper. Chris Tar and Mario Guajardo-Céspedes for the early experiment discussions and feedback. Christian Frank, Duc Dung Nguyen, Alex Tudor, and Dalia El Badawy for helping answer questions about AudioPaLM.
Appendix A Appendix
Input Type Before Tokenization Input Ids
Speech [English Speech] 50,210,245, $\ldots$ 240, 503, 32050, 32210, 32245, $\ldots$
Transcription [English Text] Hello World . 59, 294, 691, $\ldots$
Table 3: Example of the speech and transcript inputs given to our model. The speech input is composed of a prefix containing the language and the input modality. Text will be tokenized using the LLMs tokenizer and an offset will be applied to the audio token to match the tokens that were reserved within the audio token vocabulary. Bold numbers represent the audio tokens before tokenization and after the offset is applied to the audio tokens.
A.1 Training Setup
Ni et al. (2022) showed that applying a contrastive loss to sentence encoders leads to improved retrieval performance in downstream tasks. After initializing our model from the PaLM 2, we use a contrastive loss Hadsell et al. (2006).
$L=-\frac{1}{N}\sum_{i=1}^{N}\frac{e^{\text{sim}(\bm{x}_{i},\bm{y}_{i})}}{\sum_{j=1}^{N}e^{\text{sim}(\bm{x}_{i},\bm{y}_{j})}}$ (1)
Using equation 1, our multi-modal DE will learn from paired speech and text embeddings $(\bm{x}_{i},\bm{y}_{i})$, where $\bm{y}_{i}$ is considered as a positive example to $\bm{x}_{i}$ while all other examples where $i\neq j$ are negative ones. The model should learn to bring the positive transcriptions closer to the corresponding speech sample, while pushing away all the other negative transcriptions. In our training, the positive and negative distinction is done within the training batch. Hence, we apply an in-batch softmax as part of our loss computation. Lastly, sim() is a similarity function formulated as the dot product between the speech sample and the transcription embeddings.
To train our model, we use the sum of a contrastive loss with a spreadout loss Zhang et al. (2017) of both the speech and text embeddings. We calculate the contrastive loss Yang et al. (2019) in a bidirectional way, by adding the loss in the speech-to-text and the text-to-speech direction.
We use the Adam Kingma and Ba (2014) optimizer with a learning rate of $1.0\times 10^{-3}$ with linear ramp cosine decay scheduler with 2.5k warm up steps. We use a dropout probability of $0.1$ and train for 100k steps with a batch size of 1024.
A.2 Expressing Tasks
For training and inference, we found that using a prefix improves speech-to-text retrieval performance. Therefore, we pre-pend a prefix containing the language and modality shown in in Table 3. In the case of a speech utterance, the prefix will be tokenized with the LLMs tokenizer and the remaining will be converted to audio tokens.
A.3 Data
Dataset Type Task Langs. Split Size
CoVoST-2 Speech S2T 21 Train 900 h.
FLEURS Speech S2T 102 Test 283 h.
FLEURS Speech S2TT 102 Test 283 h.
Wikimatrix Text MT 4 Train 9M sents.
Table 4: Training and evaluation datasets. CoVoST-2 is used for speech-to-text retrieval (S2T), Wikimatrix is for machine translation retrieval (MT), and FLEURS is for evaluating X $\to$ En speech-to-text translation retrieval (S2TT) and also speech-to-text retrieval (S2T).
# Sents. X $\to$ En
German (de) 6.2M
Polish (pl) 2.1M
French (fr) 705k
Dutch (nl) 570k
Table 5: Number of parallel sentences used in the machine translation mixture from Wikimatrix corpus.
Table 4 shows the training and evaluation datasets we used through out our experiments. We used 21 languages CoVoST-2 to train our model on speech-to-text retrieval which amounts to approximately 900 hours of speech. To evaluate our models speech-to-text retrieval capabilities, we evaluate on FLEURS speech-to-text test split on 102 languages. We use FLEURS speech-to-text translation test split to evaluate our models abilities on tasks that require cross-lingual and cross-modal knowledge. We evaluate of 4 different languages: German, Polish, French, and Dutch.
We find that combining speech-to-text retrieval data and readily available translation data improves our models cross-lingual and cross-modal abilities. Table 5 shows the number of parallel sentences we used during training from X $\to$ En.

### v2

@@
-Transforming LLMs into Cross-modal and Cross-lingual Retrieval Systems
+Using LLMs as Speech-to-Text Retrieval Systems
 Abstract
@@
 2 Method
-We train a transformer-based DE model that encodes speech and text given a dataset, $\emph{D}=\{(x_{i},y_{i})\}$, where $x_{i}$ is a speech utterance and $y_{i}$ is its transcription.
+We train a transformer-based DE model that encodes speech and text given a dataset $\emph{D}=\{(x_{i},y_{i})\}$, where $x_{i}$ is a speech utterance and $y_{i}$ is its transcription.
 We denote the speech and text embeddings as $\bm{x_{i}}=E(x_{i})$ and $\bm{y_{i}}=E(y_{i})$, respectively, where $E$ is a transformer-based DE that encodes speech and text.
@@
 Appendix A.1 includes more details on our training setup.
+We will refer to our proposed model as PaLM 2 DE.
 5 Experiments
 R@1 $\uparrow$ WER $\downarrow$
-PaLM 2 XXS 86.15 13.85
-mSLAM 76.9 14.6
-Table 1: PaLM 2 results for $R@1$ and WER compared against the mSLAM on 102 languages from FLEURS for speech-to-text retrieval (S2T).
+mSLAM DE Conneau et al. (2023) 76.9 14.6
+PaLM 2 DE (Proposed Model) 86.15 13.85
+Table 1: PaLM 2 DE results for R@1 and WER compared against the mSLAM DE on 102 languages from FLEURS for speech-to-text retrieval (S2T).
 We train our DE model to perform S2T, where the task is to retrieve the corresponding transcription given a speech sample.
@@
 Table 1 shows the average $R@1$ and WER for S2T for 102 languages from FLEURS.
-We compare against the mSLAM model from Conneau et al.
+We compare against the mSLAM DE model from Conneau et al.
 (2023), a model trained on 426k hours of S2T data in 51 languages and fine-tuned on FLEURS training data.
-Our model significantly outperforms the mSLAM baseline in $R@1$ and WER metrics despite being trained with only 1/10 of the data and having been initialized from a text-only LLM.
+Our model significantly outperforms the mSLAM DE baseline in $R@1$ and $WER$ metrics despite being trained with only 1/10 of the data and having been initialized from a text-only LLM.
 More importantly, our model was only trained on the 21 languages in CoVoST-2 and never fine-tuned on the FLEURS training data.
@@
 R@1 $\uparrow$
-Language Group (#) PaLM 2 (%) mSLAM (%)
-Afro-Asiatic (7) 90.82 73.67
-Atlantic-Congo (15) 79.47 86.77
-Austro-Asiatic (3) 41.71 47.90
-Austronesian (6) 85.74 75.50
-Dravidian (4) 90.46 65.70
-Indo-European (51) 92.38 84.62
-Japonic (1) 63.23 5.80
-Kartvelian (1) 74.57 70.50
-Koreanic (1) 45.81 5.20
-Kra-Dai (1) 37.12 3.20
-Mongolic (1) 98.10 70.70
-Nilo-Saharan (1) 94.53 91.00
-Sino-Tibetan (3) 61.91 3.40
-Turkic (4) 91.89 81.28
-Uralic (3) 93.38 91.40
+Language Group (#) mSLAM DE PaLM 2 DE
+Conneau et al. (2023) (Proposed Model)
+Afro-Asiatic (7) 73.67 90.82
+Atlantic-Congo (15) 86.77 79.47
+Austro-Asiatic (3) 47.90 41.71
+Austronesian (6) 75.50 85.74
+Dravidian (4) 65.70 90.46
+Indo-European (51) 84.62 92.38
+Japonic (1) 5.80 63.23
+Kartvelian (1) 70.50 74.57
+Koreanic (1) 5.20 45.81
+Kra-Dai (1) 3.20 37.12
+Mongolic (1) 70.70 98.10
+Nilo-Saharan (1) 91.00 94.53
+Sino-Tibetan (3) 3.40 61.91
+Turkic (4) 81.28 91.89
+Uralic (3) 91.40 93.38
 Table 2: FLEURS S2T (R@1) performance broken down by language groups. Bold represents better performance. Numbers in parenthesis represent the number of languages within the language group.
 Table 2 shows the $R@1$ language group breakdown for S2T on FLEURS.
-We find that although we only trained on 21 languages, our model significantly outperforms mSLAM in 13 of the 15 language groups.
+We find that although we only trained on 21 languages, our model significantly outperforms mSLAM DE in 13 of the 15 language groups.
 These results are consistent with the experiments in Hassid et al.
@@
 5.2 Evaluating on Cross-Modal and Cross-Lingual Tasks
-Figure 3: BLEU scores for FLEURS zero-shot S2TT when training on Transcripts or Transcripts + Translations. Combining transcripts and translation data improves zero-shot S2TT retrieval.
+Figure 3: BLEU scores for FLEURS zero-shot S2TT when training on Transcripts or Transcripts + Translations for PaLM 2 DE. Combining transcripts and translation data improves zero-shot S2TT retrieval.
 We evaluate on S2TT to gauge the cross-modal and cross-lingual capabilities of our model.
@@
 We would like to thank Shankar Kumar and Ankur Bapna for the valuable feedback on the draft of the paper.
-Chris Tar and Mario Guajardo-Céspedes for the early experiment discussions and feedback.
+Chris Tar, Mario Guajardo-Céspedes, and Jason Riesa for the early experiment discussions and feedback.
 Christian Frank, Duc Dung Nguyen, Alex Tudor, and Dalia El Badawy for helping answer questions about AudioPaLM.

### v3

@@
 Additionally, our model demonstrates cross-lingual speech and text matching, which is further enhanced by readily available machine translation data.
+[…]
+[…]
 Transforming LLMs into Cross-modal and Cross-lingual Retrieval Systems
@@
 (2023); Hassid et al.
+(2023); Gong et al.
+(2023); Peng et al.
 (2023).
@@
     We show that our model exhibits cross-lingual speech and text matching without training on this type of data.
-At the same time, we find that cross-lingual speech and text matching is furthered improved by training on readily available machine translation data.
+At the same time, we find that cross-lingual speech and text matching is further improved by training on readily available machine translation data.
 2 Method
@@
 (2023) as the speech encoder and take the middle layer as the embedding for $x_{i}$.
-Additionally, we generate audio tokens at 25Hz using k-means clustering, resulting in a set of 1024 possible audio tokens.
+Additionally, we generate audio tokens at 25Hz using k-means clustering  [^1] We use the USM-v2 audio tokenizer from Rubenstein et al.
+(2023).
 We will refer to this as our audio token vocabulary.
@@
 mSLAM DE Conneau et al. (2023) 76.9 14.6
-PaLM 2 DE (Proposed Model) 86.15 13.85
+PaLM 2 DE (Proposed Model) 86.7 13.4
 Table 1: PaLM 2 DE results for R@1 and WER compared against the mSLAM DE on 102 languages from FLEURS for speech-to-text retrieval (S2T).
@@
 In Figure 2 we break down the R@1 scores based on seen and unseen languages during training.
-We find that our model performs best on the 20 languages that are within the training and evaluation data, but still perform significantly well on the remaining 82 unseen languages.
+We find that our model performs best on the 20 languages that are within the training and evaluation data, but still perform well on the remaining 82 unseen languages.
 We hypothesize this is due to the vast textual multilingual data our backbone LLM has seen during pre-training.
@@
 R@1 $\uparrow$
-Language Group (#) mSLAM DE PaLM 2 DE
+Language Group (#) mSLAM DE PaLM 2 DE ##\## Wins
 Conneau et al. (2023) (Proposed Model)
-Afro-Asiatic (7) 73.67 90.82
-Atlantic-Congo (15) 86.77 79.47
-Austro-Asiatic (3) 47.90 41.71
-Austronesian (6) 75.50 85.74
-Dravidian (4) 65.70 90.46
-Indo-European (51) 84.62 92.38
-Japonic (1) 5.80 63.23
-Kartvelian (1) 70.50 74.57
-Koreanic (1) 5.20 45.81
-Kra-Dai (1) 3.20 37.12
-Mongolic (1) 70.70 98.10
-Nilo-Saharan (1) 91.00 94.53
-Sino-Tibetan (3) 3.40 61.91
-Turkic (4) 81.28 91.89
-Uralic (3) 91.40 93.38
-Table 2: FLEURS S2T (R@1) performance broken down by language groups. Bold represents better performance. Numbers in parenthesis represent the number of languages within the language group.
+Afro-Asiatic (7) 73.67 84.22 5
+Atlantic-Congo (14) 86.77 70.41 1
+Austro-Asiatic (2) 47.90 34.42 0
+Austronesian (6) 75.50 90.73 6
+Dravidian (4) 65.70 92.06 4
+Indo-European (51) 84.62 95.32 49
+Japonic (1) 5.80 91.54 1
+Kartvelian (1) 70.50 82.92 1
+Koreanic (1) 5.20 52.36 1
+Kra-Dai (2) 3.20 22.09 1
+Mongolic (1) 70.70 99.89 1
+Nilo-Saharan (1) 91.00 92.52 1
+Sino-Tibetan (3) 3.40 90.66 3
+Turkic (5) 81.28 92.86 4
+Uralic (3) 91.40 99.04 3
+All (102) 76.90 86.72 81
+Table 2: FLEURS S2T (R@1) performance by language groups.
+Bold represents better performance.
+Numbers in parenthesis are the number of languages within the language group.
+##\## Wins is the number of languages where PaLM 2 DE outperforms mSLAM in the language group.
 Table 2 shows the R@1 language group breakdown for S2T on FLEURS.
@@
 Figure 3: BLEU scores for FLEURS zero-shot S2TT when training on Transcripts or Transcripts + Translations for PaLM 2 DE. Combining transcripts and translation data improves zero-shot S2TT retrieval.
-We evaluate on S2TT to gauge the cross-modal and cross-lingual capabilities of our model.
-We show that we can further improve S2TT by simply training on a mixture of S2T and translation data without using any S2TT training data.
+We evaluate on S2TT to gauge the cross-modal and cross-lingual capabilities of our model. We show we can improve S2TT by simply combining S2T and translation data without S2TT training data.
 5.2.1 Zero-Shot S2TT
@@
 (2023) demonstrated that pre-trained LLMs can be adapted to various tasks that required text and speech understanding.
-On the other hand, several works aim to build joint speech and text representations.
+On the other hand, several works aim to build joint speech and text representations Khurana et al.
+(2022); Gow-Smith et al.
+(2023).
 Chung et al.
@@
 8 Acknowledgements
-We would like to thank Shankar Kumar and Ankur Bapna for the valuable feedback on the draft of the paper.
+We would like to thank Shankar Kumar, Ankur Bapna, and the anonymous reviewers for the valuable feedback on the draft of the paper.
 Chris Tar, Mario Guajardo-Céspedes, and Jason Riesa for the early experiment discussions and feedback.
@@
 Table 5 shows the number of parallel sentences we used during training from X $\to$ En.
+A.4 Performance Breakdown By Language
+Table 6 includes the PaLM 2 DE R@1 for each language found in FLEURS. We also include the language group from Table 2 and the number of examples found within each S2T test set.
+Idx Language Name Code Family ##\## Examples R@1
+mSLAM PaLM 2 DE
+1 Afrikaans af Indo-European 414 90.1 99.3
+2 Amharic am Afro-Asiatic 516 34.1 69.6
+3 Arabic ar Afro-Asiatic 427 82.7 98.8
+4 Armenian hy Indo-European 929 50.3 89.7
+5 Assamese as Indo-European 980 81.5 87.4
+6 Asturian ast Indo-European 946 90.1 100.0
+7 Azerbaijani az Turkic 918 83.0 98.4
+8 Belarusian be Indo-European 955 90.2 97.2
+9 Bengali bn Indo-European 911 83.5 84.6
+10 Bosnian bs Indo-European 923 95.5 99.8
+11 Bulgarian bg Indo-European 657 95.1 100.0
+12 Burmese my Sino-Tibetan 870 2.4 19.3
+13 Cantonese yue Sino-Tibetan 819 2.4 83.6
+14 Catalan ca Indo-European 938 93.2 100.0
+15 Cebuano ceb Austronesian 532 79.8 94.9
+16 Croatian hr Indo-European 914 98.0 99.8
+17 Czech cs Indo-European 720 98.1 99.6
+18 Danish da Indo-European 929 94.1 99.9
+19 Dutch nl Indo-European 364 95.3 100.0
+20 English en Indo-European 647 96.0 99.1
+21 Estonian et Uralic 892 95.6 99.9
+22 Filipino fil Austronesian 928 73.1 89.1
+23 Finnish fi Uralic 916 93.0 98.9
+24 French fr Indo-European 675 90.7 100.0
+25 Fula ff Atlantic-Congo 649 81.4 81.7
+26 Galician gl Indo-European 927 90.9 100.0
+27 Ganda lg Atlantic-Congo 705 90.7 75.7
+28 Georgian ka Kartvelian 978 70.5 82.9
+29 German de Indo-European 841 91.2 100.0
+30 Greek el Indo-European 649 81.2 73.2
+31 Gujarati gu Indo-European 1000 77.0 95.9
+32 Hausa ha Afro-Asiatic 557 84.5 83.1
+33 Hebrew he Afro-Asiatic 792 64.0 76.0
+34 Hindi hi Indo-European 417 78.0 83.7
+35 Hungarian hu Uralic 902 85.3 98.3
+36 Icelandic is Indo-European 46 71.7 97.8
+37 Igbo ig Atlantic-Congo 869 85.8 64.9
+38 Indonesian id Austronesian 684 79.6 99.4
+39 Irish ga Indo-European 829 55.1 69.5
+40 Italian it Indo-European 857 93.5 100.0
+41 Japanese ja Japonic 650 5.8 91.5
+42 Javanese jv Austronesian 722 78.0 97.0
+43 Kabuverdianu kea Indo-European 859 95.4 99.9
+Idx Language Name Code Family ##\## Examples R@1
+mSLAM PaLM 2 DE
+44 Kamba kam Atlantic-Congo 798 89.7 81.5
+45 Kannada kn Dravidian 831 69.0 88.8
+46 Kazakh kk Turkic 841 88.7 83.1
+47 Khmer km Austro-Asiatic 765 42.1 20.3
+48 Korean ko Koreanic 382 5.2 52.4
+49 Kyrgyz ky Turkic 974 84.3 88.6
+50 Lao lo Kra-Dai 399 37.0 23.3
+51 Latvian lv Indo-European 848 97.4 100.0
+52 Lingala ln Atlantic-Congo 440 91.2 76.4
+53 Lithuanian lt Indo-European 985 96.8 98.2
+54 Luo luo Nilo-Saharan 254 91.0 92.5
+55 Luxembourgish lb Indo-European 929 80.5 74.6
+56 Macedonian mk Indo-European 967 96.1 98.8
+57 Malay ms Austronesian 749 77.7 98.7
+58 Malayalam ml Dravidian 944 62.3 88.3
+59 Maltese mt Afro-Asiatic 918 92.7 76.0
+60 Mandarin cmn Sino-Tibetan 944 5.4 100.0
+61 Maori mi Austronesian 890 64.7 65.3
+62 Marathi mr Indo-European 1005 69.8 82.4
+63 Mongolian mn Mongolic 949 70.7 99.9
+64 Nepali ne Indo-European 724 66.1 89.6
+65 Northern-Sotho nso Atlantic-Congo 738 80.8 70.3
+66 Norwegian nb Indo-European 357 91.9 100.0
+67 Nyanja ny Atlantic-Congo 745 85.5 63.6
+68 Occitan oc Indo-European 968 77.4 99.4
+69 Oriya or Indo-European 875 15.7 95.1
+70 Oromo om Afro-Asiatic 41 92.7 100.0
+71 Pashto ps Indo-European 510 84.8 91.0
+72 Persian fa Indo-European 858 85.4 100.0
+73 Polish pl Indo-European 758 95.8 99.3
+74 Portuguese pt Indo-European 914 91.9 99.9
+75 Punjabi pa Indo-European 574 70.6 96.7
+76 Romanian ro Indo-European 882 92.0 100.0
+77 Russian ru Indo-European 774 93.2 100.0
+78 Serbian sr Indo-European 700 97.7 99.1
+79 Shona sn Atlantic-Congo 920 84.1 53.9
+80 Sindhi sd Indo-European 977 71.8 85.4
+81 Slovak sk Indo-European 791 97.6 99.5
+82 Slovenian sl Indo-European 834 97.4 100.0
+83 Somali so Afro-Asiatic 1007 68.7 86.0
+84 Sorani-Kurdish ckb Indo-European 918 80.8 96.7
+85 Spanish es Indo-European 907 69.6 100.0
+86 Swahili sw Atlantic-Congo 487 91.2 86.2
+Idx Language Name Code Family ##\## Examples R@1
+mSLAM PaLM 2 DE
+87 Swedish sv Indo-European 758 94.2 100.0
+88 Tajik tg Indo-European 590 76.3 92.7
+89 Tamil ta Dravidian 582 58.0 98.1
+90 Telugu te Dravidian 471 73.5 93.0
+91 Thai th Kra-Dai 1011 3.2 20.9
+92 Turkish tr Turkic 742 84.5 100.0
+93 Ukrainian uk Indo-European 750 93.5 99.3
+94 Umbundu umb Atlantic-Congo 264 77.3 62.1
+95 Urdu ur Indo-European 299 70.6 91.3
+96 Uzbek uz Turkic 861 67.6 94.2
+97 Vietnamese vi Austro-Asiatic 850 64.5 48.6
+98 Welsh cy Indo-European 1002 82.3 96.1
+99 Wolof wo Atlantic-Congo 351 90.6 87.5
+100 Xhosa xh Atlantic-Congo 1034 90.9 30.2
+101 Yoruba yo Atlantic-Congo 816 92.4 84.6
+102 Zulu zu Atlantic-Congo 822 85.5 67.2
+All (102) 76.9 86.7
+Table 6: Language name, code, family, and number of examples for each test set found in FLEURS. We report R@1 for mSLAM and PaLM 2 DE.